OpenStudy (anonymous):

Use point by point plotting to sketch the graph of this function f(x)=1-x?

OpenStudy (mertsj):

Choose some values for x and calculate the corresponding function values and plot the points.

OpenStudy (anonymous):

F(x)=-x + 1 (I just re-wrote the equation easier for reading) y=-x + 1 F(x)=y same thing. substitute any point into the equation to get another point and plot it.

OpenStudy (anonymous):

Hint: if there are any x-intercepts or y-intercepts, u should plot them because they are the most important.

OpenStudy (anonymous):